# PixFoundation: Are We Heading in the Right Direction with Pixel-level Vision Foundation Models?

### v2

@@ Appendix A Additional implementation details
Additionally, we provide details on the SAM model that is used in the three baselines and upper bounds in our benchmarks, where we use the ViT-H variant. Finally, we provide an illustrative example of our automatic selection mechanism with the corresponding predictions on PixMMVP using Llava 1.5 (7B) in Fig. 11. Our automatic selection goes through an iterative process of prompting the selected MLLM, in our case GPT-4o, with groups of images highlighting the predicted segmentation to select the best within each group. In the final stage the best images are used to prompt the MLLM to select the final selected mask that best describes the object of interest.
| Model Name | Model Checkpoint |
| LISA | xinlai/LISA-7B-v1-explanatory |
| GLAMM | MBZUAI/GLaMM-FullScope |
| GLAMM-RegCap | MBZUAI/GLaMM-RegCap-RefCOCOg |
| Llava-G | Haozhangcx/llava_grounding_gd_vp |
| Llava 1.5 (7B) | liuhaotian/llava-v1.5-7b |
| Llava 1.5 (13B) | liuhaotian/llava-v1.5-13b |
| Cambrian-1 (8B) | nyu-visionx/cambrian-8b |
Evaluation. We also provide the details on computing the visual question answering accuracy using GPT-4o in the first protocol [19]. We use the following prompt: “Given the following question QUESTION, the correct answer is ANSWER. Does the following answer correctly answers the question, answer: RESPONSE? Respond with a Yes/No”.
Appendix B Additional qualitative analysis
@@
B.1 Baselines ablation
We show the qualitative ablation among the two baselines and upper bound using the best base MLLM Cambrian-1 (8B) in Fig. 12 on PixMMVP. The three confirm that there is grounding emerging in MLLMs that were not trained with pixel-level grounding supervision. Nonetheless, it shows that identifying when that grounding emerges is equally important in retrieving the best segmentation of the referring expression. The first baseline, attend and segment, assumes the alignment between the attention map that can be mined for the segmentation mask and the noun phrase that has the highest correspondence to the ground-truth category or noun phrase. Our findings quantitatively and qualitatively show otherwise, where grounding can emerge in different output tokens. It also shows the oracle upper bound for mask selection, PixFoundation, exhibiting better segmentation than the attend and segment, confirming on the aforementioned finding. Additionally, it shows that our simple automatic mechanism, PixFoundation, surpasses the attend and segment as well on PixMMVP.
Attend and Segment [2] (Cambrian-1 (8B))
PixFoundation (Cambrian-1 (8B))
PixFoundation (Cambrian-1 (8B))
[5pt]
@@ B.1 Baselines ablation
B.2 Attention maps visualization
In this section, we visualize the normalized attention maps, , in Fig. 13. We show two examples for Cambrian-1 (8B) from PixMMVP using the first probing where we directly prompt the model with question and options. The first row shows outstanding ability to visually ground the different noun phrases from the output text. The full output text of the first row example is: “The image provided is a cake designed to resemble a minion from the D̈espicable Mef̈ranchise. It is not a living creature and therefore cannot smile or have a tongue out. The cake is an inanimate object, crafted to mimic the appearance of a minion, which is a fictional character from the animated movie series. The design elements such as the yellow skin, blue overalls, and goggles are characteristic of the minions’ appearance in the films.” The visualization shows how the maximally attended locations for the last three noun phrases correspond to the correct locations in the image.
@@ B.2 Attention maps visualization
B.3 When does grounding emerge?
We show additional examples on when grounding emerges in multi-modal large language models, specifically in the Llava 1.5 (7B) variant, using the second probing to prompt the model to segment what is in the referring expression. Figures 14, 16, 15 and 17 show the corresponding predicted masks for the grounding that emerged highlighted in red with the maximum attention point as a black dot. Figure 8 in the main text shows the aforementioned four examples with the input prompt, the concept category and the noun phrase corresponding to the best grounding using the oracle selection and the full output text. It clearly shows that the correct output token can correspond to location or color but not necessarily the ground-truth referring expression. While some of the noun phrases and their masks from the point prompting of SAM do correspond to what the noun phrase is describing. It is not always the case, example in Fig. 16 “the flame” was not able to highlight the correct object yet it appeared in the noun phrase corresponding to the location “the top”. It also shows that the majority of correct grounding occurs near the end of the output text after the model reasoned for an answer to the question.
Additionally, this type of mining for attention maps can show potential mistakes that MLLMs fall into while performing the grounding to reason on what is in the image. An example on that is in Fig. 17, where the model potentially mistakes a crack in the wood for the minute hand. Thus, our PixFoundation serves as an interesting tool to interpret and understand how MLLMs work and reason to produce the final output with the oracle selection as an upper bound. Yet, there is an open question on how to identify the real mistake from the MLLM vs. simple misalignment between the noun phrase and the grounding emerging, since grounding could emerge with phrases describing it or parts of it instead of the object’s noun phrase itself. We leave this for future work.
@@ B.3 When does grounding emerge?
B.4 PixMMVP benchmark
Figure 18 shows additional results on PixMMVP benchmark comparing different pixel-level MLLMs with our oracle baseline using Llava 1.5 (7B). While GLAMM shows strong pixel-level visual grounding yet we have shown earlier that it is almost incapable of visual question answering which renders the model weak for general purpose tasks. On the other hand, OMG-Llava shows a better balance in pixel-level visual grounding and visual question answering as previously detailed. Nonetheless, the simple mining of attention maps from Llava 1.5 (7B) using the oracle selection which we call PixFoundation shows the strongest capability in both grounding and VQA. In fact, certain MLLMs that were trained with pixel-level visual grounding, such as LISA, have degraded the performance with respect to the hidden information already existing in powerful MLLMs that were not trained with such supervision.
@@ B.4 PixMMVP benchmark
B.5 PixCV-Bench benchmark
Figure 19 shows qualitative results on PixCV-Bench. It shows that pixel-level MLLMs struggle with segmenting the object annotated by the red box unlike our oracle baseline, PixFoundation. Indeed the attention maps from these MLLMs are looking at the right object annotated by the red box without receiving any pixel-level grounding supervision during training.
@@ B.5 PixCV-Bench benchmark
Appendix C Failure case analysis
In this section, we conduct additional failure case analysis of pixel-level MLLMs and our baselines qualitatively and quantitatively.
C.1 Failures in visual question answering
@@
C.2 Failures in pixel-level visual grounding
@@
Another failure occurring such as in the second row stems from ambiguity in the referring expression itself or failures from SAM identifying the separation between the wall and the ceiling. Hence, the oracle upper bound is generally inheriting SAM failures. However, its main purpose of showing that the hidden information within powerful MLLMs is sufficient to perform pixel-level grounding is achieved, and even surpassing pixel-level MLLMs without degrading their VQA abilities.
Appendix D Additional quantitative analysis
D.1 When grounding emerges - CV-Bench
In Fig. 23(a) we show the analysis on when grounding emerges on PixCV-Bench in terms of the frequency of the grounding location. It is worth noting that PixMMVP is more challenging than PixCV-Bench evidently from the reported IoU and accuracy metrics on both reported in Table 1. It seems on the less challenging dataset PixCV-Bench, grounding tends to emerge frequently near the beginning of the output. This might relate to PixMMVP being more challenging in terms of the level of reasoning than CV-Bench or the fact that PixMMVP poses a harder referring segmentation task than PixCV-Bench, which is mostly using the class names. However, the consistent finding among both datasets is that grounding can emerge coinciding to various concept categories whether location, color or state as shown in Fig. 23(b).
D.2 Analysis on the output length
In this section, we provide additional analysis on the output length on average through PixMMVP dataset using the first and second probing schemes. Specifically, we report the output length as the number of characters in the output, and the number of noun phrases extracted from it. The reason to study this, since it has relation to the number of noun phrases and consequently the number of masks our baselines are selecting among. Table 3 shows the average output length computed across PixMMVP dataset, comparing the three base MLLMs. We notice that Cambrian-1 (8B) generates longer outputs with a considerable margin than Llava variants. Hence, we believe the superiority of the oracle upper bound with Cambrian-1 in the grounding has strong correlation to producing longer outputs with more attention maps to mine and select from, than Llava variants.
| Model Name | Probing | Output Length | # Noun Phrases |
| Llava 1.5 (7B) | First | 44.2 | 2.3 |
| Llava 1.5 (13B) | First | 45.3 | 2.4 |
| Cambrian-1 (8B) | First | 313.8 | 15.2 |
| Llava 1.5 (7B) | Second | 92.6 | 5.2 |
| Llava 1.5 (13B) | Second | 97.2 | 5.5 |
| Cambrian-1 (8B) | Second | 561.3 | 27.3 |

### v4

@@ Appendix A Additional implementation details
Additionally, we provide details on the SAM model that is used in our baselines and interpretability tool, where we use the ViT-H variant. For the automatic and the oracle variants of PixFoundation with Qwen2.5-VL, we use the output point generated directly from the model instead of mining the attention maps, where we provide these results for reference. Nonetheless, we maintain the mask selection among the three masks corresponding to each point. Finally, we provide an illustrative example of our automatic selection mechanism with the corresponding predictions on PixMMVP using LLaVA 1.5 (7B) in Fig. 7. Our automatic selection goes through an iterative process of prompting the selected MLLM, in our case GPT-5.1, with N images highlighted with the predicted segmentation to select the best within each group. In the final stage, the best images are used to prompt the MLLM to select the final mask that best describes the object of interest.
keyboard (annotated by the red box)
bottle
kite
drinking glass
Evaluation. We also provide the details on computing the visual question answering accuracy using GPT-4o in the first probing [24] (i.e., ), or when evaluating the language prompt sensitivity in the third probing with GPT-4o (i.e., w/ GPT). We use the following prompt: “Given the following question QUESTION, the correct answer is ANSWER. Does the following answer correctly answer the question, answer: RESPONSE? Respond with a Yes/No”. As for the mean intersection over union on PixMMVP (i.e., ), we provide two results: the results on all images, including the ones that do not have an object of interest and their referring expression is marked as “None”, and when excluding this set of images. The remaining results in the supplementary for the mean intersection over union are excluding this set unless stated otherwise. Note that our benchmarking was conducted on an A6000 48G GPU-equipped machine.
Prompt sensitivity. For the language prompt sensitivity, we generate eight random locations in both the question and the choices to change, excluding the enumeration of the choices. Similarly, we randomly select eight locations in the instruction prompting the model to generate an option’s letter. Table 5 and Table 6 show the prompt templates that are used in the language prompt sensitivity evaluation in the VQA and the visual grounding, respectively.
| # | Prompt Template |
[…]
| 3 | QUES CHOICES INSTRUCT |
| 4 | Question-QUES CHOICES. INSTRUCT. Answer- |
[…]
| 6 | Question, QUES CHOICES. INSTRUCT. Answer, |
[…]
| 9 | QUES CHOICES INSTRUCT |
[…]
| # | Prompt Template |
| 1 | Can you please segment/detect EXPR in the given image |
| 2 | Can you segment/detect EXPR in this image? |
| 3 | Can you identify EXPR in this image? (with grounding) |
| 4 | Identify EXPR in the scene, with grounding. |
| 5 | Locate the EXPR and output a tight mask/box. If the object does not exist in the image don’t generate any masks/boxes. |
| 6 | Output mask/box for the EXPR |
Appendix B Additional quantitative analysis
B.1 PixFoundation using open-source models
In our automatic baseline, we replace GPT-5.1, which is a closed-source model, with another open-source model, in our case, Cambrian (8B). Table 7 shows the results on PixMMVP for PixFoundation, the automatic baseline that still surpasses the best pixel-level MLLM, OMG-LLaVA, that was before the release of our benchmark. More importantly, this baseline confirms that even with the use of a self-contained model such as Cambrian, without additional help from GPT-5.1, it can still compete with these unified pixel-level supervised models.
| Method | PixMMVP | ||||
| OMG LLaVA (7B) | 12.0 | 12.0 | 13.8 | 38.8 | 18.3 |
| LLaVA 1.5 (7B) + (a+s) | 27.3 | 28.0 | 7.8 | 14.1 | 18.8 |
| LLaVA 1.5 (13B) + (a+s) | 39.3 | 30 | 8.5 | 13.9 | 20.5 |
| Cambrian (8B) + (a+s) | 52.0 | 52.0 | 15.5 | 13.3 | 23.9 |
| Cambrian (8B) + PixFoundation (Ours) | 52.0 | 52.0 | 44.7 | 41.4 | 48.1 |
| Cambrian (8B) + PixFoundation (Ours) | 52.0 | 52.0 | 24.6 | 25.8 | 34.5 |
B.2 Random baseline
Our baselines rely on the maximum attention per output noun phrase to prompt SAM for the segmentation mask. Nonetheless, as a lower bound analysis, we evaluate the performance if we use a random point as a prompt instead. For fair comparison, we generate random points with the count of output masks that the oracle has to select among (i.e., the number of the output noun phrases). We conduct this ablation on PixMMVP using LLaVA 1.5 (7B) base MLLM, with random point prompts followed by the oracle selection among their SAM masks. Figure 9 shows that random + oracle lags behind the correct one using the maximum point (i.e., 1st) with around 28.9%. More importantly, we confirm the stability of the results if we select the second-best or third maximum attention (i.e., 2nd or 3rd), which are on par with the maximum attention (i.e., 1st).
B.3 Conventional visual grounding benchmarks
We evaluate our upper bound on the conventional benchmarks for RefCOCO/RefCOCO+/RefCOCOg [10, 17] with respect to state-of-the-art methods in pixel-level MLLMs prior to the release of our benchmark. Table 8 shows that our upper bound that relies on extracting the output noun phrases, mining their attention maps and selecting the best mask output in relation to these attention maps is competitive to state-of-the-art methods. However, the main goal of PixFoundation is to inspect when visual grounding occurs in MLLMs with respect to the output token. We still show the results on refCOCO benchmarks for reference. Note that unlike refCOCO, our benchmarks pose a greater challenge for pixel-level MLLMs since they are focused on vision-centric tasks that require grounding to answer the questions. Furthermore, our paired evaluation enables the evaluation of the grounding in relation to the VQA task to better study these MLLMs. Finally, unlike refCOCO variants, our PixMMVP especially provides an out-of-distribution challenging evaluation benchmark, where the majority of these MLLMs have been trained on refCOCO training data.
| Method | refCOCO | refCOCO+ | refCOCOg | |||||
| val | testA | testB | val | testA | testB | val | test | |
| LISA | 74.1 | 76.5 | 71.1 | 62.4 | 67.4 | 56.5 | 66.4 | 68.5 |
| LISA (ft) | 74.9 | 79.1 | 72.3 | 65.1 | 70.8 | 58.1 | 67.9 | 70.6 |
| OMG LLaVA (7B) | 75.6 | 77.7 | 71.2 | 65.6 | 69.7 | 58.9 | 70.7 | 70.2 |
| OMG LLaVA (7B) (ft) | 78.0 | 80.3 | 74.1 | 69.1 | 73.1 | 63.0 | 72.9 | 72.9 |
| Cambrian (8B) + PixFoundation | 77.8 | 81.1 | 72.8 | 73.2 | 79.5 | 70.1 | 75.5 | 77.5 |
B.4 When grounding emerges - PixCV-Bench
In Fig. 10(a), we show the analysis of when grounding emerges on PixCV-Bench in terms of the location in the output. It is worth noting that PixMMVP is more challenging than PixCV-Bench, evidently from the reported IoU and accuracy metrics on both with respect to Table 1 and Table 2 main submission. It seems on the less challenging dataset PixCV-Bench, grounding tends to emerge frequently near the beginning of the output. This might relate to PixMMVP being more challenging in terms of the level of reasoning than PixCV-Bench or the fact that PixMMVP poses a harder referring segmentation task than PixCV-Bench, which mostly uses the class names. Another difference is that PixMMVP is out of the distribution of the seen datasets for most of these MLLMs. However, the consistent finding among both datasets is that grounding can emerge coinciding with various concept categories, whether location, color or state, as shown in Fig. 10(b). Note that across this analysis, we compute the frequency per object in the referred expression corresponding to the visual question. Hence, if we have two objects in one visual question, such as in the relative positioning questions, each object’s concept, corresponding to the emergence, is computed as part of our analysis.
Attend and Segment [3] (Cambrian (8B))
PixFoundation (Cambrian (8B))
PixFoundation (Cambrian (8B))
| Model Name | Prompt w/o Expr. | Referring Expression |
| RGA (7B) | 52.6 | 52.7 |
| Qwen2.5-VL (7B) | 15.6 | 28.2 |
B.5 Grounding prompt sensitivity
We evaluate the language prompt sensitivity in the visual grounding task, aside from the VQA task provided in the main submission. Table 9 compares RGA with its baseline MLLM, Qwen2.5-VL, on PixMMVP using the mean intersection over union. We conduct two variations per prompt template from Table 6, resulting in 12 variations for the prompt without the referring expression being impacted. The language variations randomly modify eight characters in the grounding prompt, where modifications can be either omission, addition or transposition. We also perform separate variations on the referred expression only within the aforementioned three modifications. Across both prompt variations, whether on the grounding prompt without the referred expression or modifying the referring expression only, RGA shows stronger results and better stability than Qwen2.5-VL. The latter exhibits higher sensitivity to the language prompt. It is mainly impacted by the prompt template, where it performs better when instructed with the “Locate” keyword than other words in the template to find the object, aside from its sensitivity to spelling errors.
Appendix C Additional qualitative analysis
@@
C.1 Baselines ablation
We show the qualitative ablation among the baseline and our interpretability mechanism using the best base MLLM Cambrian (8B) in Fig. 11 on PixMMVP. The three confirm that there is pixel-level grounding emerging in MLLMs that were not trained with mask supervision. Nonetheless, it shows that identifying when that grounding emerges is equally important in retrieving the best segmentation of the referring expression. The first baseline, attend and segment, assumes the alignment between the attention map that can be mined for the segmentation mask and the noun phrase that has the highest correspondence to the queried category or noun phrase. Our findings quantitatively and qualitatively show otherwise, where grounding can emerge in different output tokens. It also shows the oracle upper bound for mask selection, PixFoundation, exhibiting better segmentation than the attend and segment, confirming the aforementioned finding. Additionally, it shows that our simple automatic mechanism, PixFoundation, surpasses the attend and segment as well on PixMMVP.
@@ C.1 Baselines ablation
C.2 Attention maps visualization
In this section, we visualize the normalized attention maps, , in Fig. 12. We show two examples for Cambrian (8B) from PixMMVP using the first probing where we directly prompt the model with question and options. The first row shows outstanding ability to visually ground the different noun phrases from the output text. The full output text of the first row example is: “The image provided is a cake designed to resemble a minion from the D̈espicable Mef̈ranchise. It is not a living creature and therefore cannot smile or have a tongue out. The cake is an inanimate object, crafted to mimic the appearance of a minion, which is a fictional character from the animated movie series. The design elements such as the yellow skin, blue overalls, and goggles are characteristic of the minions’ appearance in the films.” The visualization shows how the maximally attended locations for the last three noun phrases correspond to the correct locations in the image.
@@ C.2 Attention maps visualization
C.3 When does grounding emerge?
We show additional examples of when grounding emerges in multi-modal large language models, specifically in the LLaVA 1.5 (7B) variant, using the second probing to prompt the model to segment what is in the referring expression. Figures 13, 14, 15 and 16 show the corresponding predicted masks for the grounding that emerged, highlighted in red with the maximum attention point as a black circle. Figure 4 main submission shows the aforementioned four examples with the referred expression, the concept category and the noun phrase corresponding to the best grounding using the oracle selection and the full output text. It clearly shows that the correct output token can correspond to location or color, but not necessarily the queried referring expression. While some of the noun phrases and their masks, from the SAM point prompting, correspond to what the noun phrase is describing. It is not always the case, for example, in Fig. 15 “the flame” was not able to highlight the correct object, yet it appeared in the noun phrase corresponding to the location “the top”. While few scenarios might have the grounding coinciding with multiple noun phrases, such as in Fig. 13, “a butterfly” and “orange wings”. Nonetheless, it is still an important insight that the segmentation can emerge corresponding to noun phrases that do not correspond to the exact referred expression. Our PixFoundation serves as an interesting tool to interpret and understand how MLLMs work and reason to produce the final output with the oracle selection as an upper bound.
In summary, we provide four evidences that pixel-level grounding can emerge corresponding to noun phrases that do not match the exact referred expression, as follows: (i) The attend and segment that rely on SpaCy embeddings lag behind our automatic and oracle mask selection, indicating that the noun phrases closest to the referred expressions are not necessarily where the optimal segmentation emerges. (ii) The point accuracy evaluation has confirmed this further, which removes any confounding factors from using three masks per noun phrase in the mask selection due to the point prompt ambiguity. The point accuracy only evaluates the emergent grounding with respect to the noun phrases. (iii) We show quantitative analysis on the location and the concept categories of the noun phrases where the grounding emerges, which confirms the previous result. (iv) We show qualitative analysis to confirm this further.
C.4 PixMMVP benchmark
Figure 17 shows additional results on PixMMVP benchmark comparing different pixel-level MLLMs with our oracle baseline using LLaVA 1.5 (7B). While GLAMM shows strong pixel-level visual grounding yet we have shown earlier that it is almost incapable of visual question answering, which renders the model weak for general-purpose tasks.
@@ C.4 PixMMVP benchmark
On the other hand, OMG-LLaVA shows a better balance in pixel-level visual grounding and visual question answering as previously detailed. Nonetheless, the simple mining of attention maps from LLaVA 1.5 (7B) using the oracle selection which we call PixFoundation shows the strongest capability in both grounding and VQA. In fact, certain MLLMs that were trained with pixel-level visual grounding, such as LISA, have degraded the performance with respect to the hidden information already existing in powerful MLLMs that were not trained with such supervision.
C.5 PixCV-Bench benchmark
Figure 18 shows qualitative results on PixCV-Bench. It shows that pixel-level MLLMs struggle with segmenting the object annotated by the red box, unlike our oracle baseline, PixFoundation. Indeed the attention maps from these MLLMs are looking at the right object annotated by the red box without receiving any pixel-level grounding supervision during training.
Appendix D Analysis on the output length
In this section, we provide additional analysis on the output length on average through PixMMVP dataset using the first and second probing schemes. Specifically, we report the output length as the number of characters in the output, and the number of noun phrases extracted from it. The reason to study this, since it has relation to the number of noun phrases and consequently the number of masks our interpretability mechanism is selecting among. Table 10 shows the average output length computed across PixMMVP dataset, comparing the three base MLLMs. We notice that Cambrian (8B) generates longer outputs with a considerable margin than LLaVA variants. Hence, we believe the superiority of the oracle upper bound with Cambrian in the grounding has strong correlation to producing longer outputs with more attention maps to mine and select from, than LLaVA variants. Nonetheless, it makes it more challenging for the automatic baseline.
| Model Name | Probing | Output Length | # Noun Phrases |
| LLaVA 1.5 (7B) | First | 44.2 | 2.3 |
| LLaVA 1.5 (13B) | First | 45.3 | 2.4 |
| Cambrian (8B) | First | 313.8 | 15.2 |
| LLaVA 1.5 (7B) | Second | 92.6 | 5.2 |
| LLaVA 1.5 (13B) | Second | 97.2 | 5.5 |
| Cambrian (8B) | Second | 561.3 | 27.3 |
@@ Appendix D Analysis on the output length
Appendix E Failure Cases Analysis
In this section, we conduct an additional failure case analysis of pixel-level MLLMs and our baselines qualitatively and quantitatively. We start with evaluating whether the failures of these MLLMs occur in visual grounding, VQA or both. Since our evaluation is paired, it enables such an analysis to have a better understanding of where the failures are stemming from. Figure 19 shows the frequency of failures per category, where the majority stem from failures in both, especially in the pixel-level MLLMs, except RGA, which shows the majority of failures are in the VQA. The standard MLLMs perform better in the VQA than grounding, as expected. However, improving pixel-level visual grounding in a naive manner that neglects the language capability eventually resulted in failures in both. On the other hand, RGA shows the highest success on both VQA and visual grounding, showing it as a promising direction, yet, as shown in the prompt sensitivity results, there are still aspects that can be improved to ensure they do not lose the language capabilities and robustness that existed in standard MLLMs. We encourage better training recipes, datasets and design choices to enable visual grounding in MLLMs without sacrificing their language abilities that are tightly coupled with grounding.
E.1 Failures in Visual Question Answering
@@
E.2 Failures in Pixel-level Visual Grounding
@@
Another failure occurring such as in the second row stems from ambiguity in the referring expression itself or failures from SAM identifying the separation between the wall and the ceiling. Hence, the oracle upper bound is generally inheriting SAM failures. However, its main purpose of showing that the hidden information within powerful MLLMs is sufficient to perform pixel-level grounding is achieved, and even surpasses some of the pixel-level MLLMs without degrading their VQA abilities.
Appendix F Impact Statement
Appendix G Limitations
Note that our baselines do entail a computational overhead with the use of the mask selection process. Nonetheless, the benefit from exploring what is already learned in these MLLMs through mining the attention maps with an understanding of when grounding emerges, provides greater benefit to interpretability. We believe interpretability of MLLMs is a crucial aspect when following a responsible approach to AI.